OpenStudy (anonymous):

how would this function appear on a graph f(x) = -3x?

terenzreignz (terenzreignz):

Functions of the form f(x) = ax + b, where a is not 0 and b is any real number, then it will always be a line.

OpenStudy (anonymous):

oh i see what about this one f(x) = 3 - x

OpenStudy (anonymous):

Its still a line, you would just rearrange it as f(x)=-x+3

OpenStudy (anonymous):

Here's graphs of those two functions

OpenStudy (anonymous):

thank you thats what i needed(: